Maple Mount Fishery is a canning company in Astoria. The company uses a normal costing system in which factory overhead is applied on the basis of direct labor costs. Budgeted factory overhead for the year was $680,000, and management budgeted $354,000 of direct labor costs. During the year, the company incurred the following actual costs. Direct materials used Direct labor Factory overhead $375,000 309,000 655,600 The January 1 balances of inventory accounts are shown below Materials-all direct Work-in-process Finished goods $61,900 45,000 27,600 The December 31 balances of these inventory accounts were ten percent lower than the balances at the beginning of the year. The cost of goods manufactured during the year is: (Round your predetermined overhead rate to 1 decimal place.)Multiple Choice $1,271100 $1,278,360. $1,316,100. $1,346,860 $1,275,600
Answer #1

Overhead rate = 680000/354000 = 1.9 of direct labor cost

Total manufacturing cost = 375000+309000+(309000*1.9) = $1271100

Cost of goods manufactured = 45000+1271100-40500 = $1275600

So answer is e) $1275600